OpenStudy (anonymous):

@LightYagami96 I have one more xD? ; 1. Examine the following drawing. Show the sum of the areas of the four smaller rectangles and the sum of the area of the large rectangle. In your own words, explain how they are related and why this is so. (I'll p

OpenStudy (anonymous):

*Ill post pic in comments (failed) ok one sec

OpenStudy (anonymous):

@johnweldon1993 can you help im stuck, i suck at math lol

OpenStudy (anonymous):

i think i have the answer but im not sure.

OpenStudy (johnweldon1993):

First, what is the area of a rectangle? \(\large A = L \times W\) right? so we have Area of rectangle A.....\(\large = 3 x\) Area of rectangle B... \(\large =21\) Area of rectangle C....\(\large= x^2\) Area of rectangle D \(\large = 7x\) Now...adding all those up, we have \(\large x^2 + 10x + 21\) Now to tackle the whole rectangle.. the total width is 3 + x and the total length is 7 + x So the total area being length times width we have \[\large Area = (3 + x)\times (7 + x)\] \[\large 21 + 10x + x^2\] Look familiar?

OpenStudy (anonymous):

lol "Now to tackle the whole rectangle.." i just learned something new lol @johnweldon1993

OpenStudy (johnweldon1993):

Haha gotta use engaging commentary when tutoring :P

OpenStudy (anonymous):

lol